OpenStudy (anonymous):

find eigen values and functions:

OpenStudy (anonymous):

\[y''+\lambda=0\]

OpenStudy (anonymous):

y'(0)=0, y(L)=0

OpenStudy (anonymous):

\[m^2=\pm\sqrt{\lambda}i\]

OpenStudy (anonymous):

m= imean

OpenStudy (anonymous):

\[y=c_1cos(\sqrt{\lambda}t)+c_2sin(\sqrt{\lambda}t)\]

OpenStudy (anonymous):

\[y'(t)=(\sqrt{\lambda})-c_1sin(\sqrt{\lambda}t)+\sqrt{\lambda}c_2cos(\sqrt{\lambda}t)\]

OpenStudy (anonymous):

y'(0)=0

OpenStudy (anonymous):

\[c_2=0\]

OpenStudy (anonymous):

since \[c_2=0\] \[y=c_1cos(\sqrt{\lambda}t)\]

OpenStudy (anonymous):

\[0=c_1cos(\sqrt{\lambda}t)\]

OpenStudy (anonymous):

only happens when c1= 0 which is a trivial solution

OpenStudy (anonymous):

or

OpenStudy (anonymous):

\[\sqrt{\lambda}L=\frac{\pi}{2}+\pi n\]

OpenStudy (anonymous):

\[\frac{\pi+2\pi n}{2}=L\sqrt{\lambda}\]

OpenStudy (anonymous):

for some reason they get 2n-1)pi

OpenStudy (anonymous):

but isn't thta the same as 2n+1)pi

OpenStudy (anonymous):

only one is starting backwards and the other is forward

OpenStudy (unklerhaukus):

\[y''+\lambda=0\]\[m^2+\lambda=0\]\[(m+i\sqrt{\lambda})(m-i\sqrt{\lambda})=0\]\[m=\pm i\sqrt\lambda\] {you already have this}

OpenStudy (unklerhaukus):

\[y_p=c_1\cos(\sqrt{\lambda}t)+c_2\sin(\sqrt{\lambda}t)\]\[y_p^\prime=-c_1\sqrt\lambda \sin(\sqrt{\lambda}t)+c_2\sqrt\lambda \cos(\sqrt{\lambda}t)\]\[y_P^{\prime\prime}=-c_1\lambda \cos\left(\sqrt{\lambda}t\right)- c_2\lambda\sin\left(\sqrt{\lambda}t\right)\]

OpenStudy (unklerhaukus):

\[y_p^\prime(t)=-c_1\sqrt\lambda \sin(\sqrt{\lambda}t)+c_2\sqrt\lambda \cos(\sqrt{\lambda}t)\]\[y_p^\prime(0)=-c_1\sqrt\lambda \sin(0)+c_2\sqrt\lambda \cos(0)=0\]\[y_p^\prime(0)=c_2\sqrt\lambda=0;\qquad\qquad c_2=0\]

OpenStudy (unklerhaukus):

\[y_p(t)=c_1\cos(\sqrt{\lambda}t)\]\[y_p(L)=c_1\cos(\sqrt{\lambda}L)=0\] \[\qquad\qquad\sqrt\lambda L=\arccos\left(0\right)\]\[\qquad\qquad\sqrt\lambda L=\frac\pi2+n\pi\]

OpenStudy (unklerhaukus):

\[\sqrt\lambda L=\frac\pi2+n\pi=\frac{\pi+2\pi n}{2}\]

OpenStudy (unklerhaukus):

\[\sqrt\lambda L=\arccos\left(0\right)=\left(n+\frac1{2}\right)\pi=\left((n+1)-\frac1{2}\right)\pi=\left(n'-\frac1{2}\right)\pi\] \[\]

OpenStudy (unklerhaukus):

your right the only difference is where n starts, this is pretty arbitrary, your answer is right

OpenStudy (unklerhaukus):

so the roots are the same as the eigen values?